OpenStudy (anonymous):

The average number of people attending a local church on Sundays dropped from 265 in October 2008 to 169 in October 2009. Find the rate of change in attendance at church.

hero (hero):

In other words, find the slope

hero (hero):

Hint: Year represents independent variable, x Attendance represents dependent variable, y

OpenStudy (anonymous):

265-169=96

hero (hero):

Will the rate of change be positive or negative?

OpenStudy (anonymous):

265-169=96 365/96=3.80

hero (hero):

Okay, now you've gone too far.

OpenStudy (anonymous):

265-169=96 365/96=.263013698 .26301369=26%

hero (hero):

Okay, now you've really lost me

hero (hero):

It didn't say find percent change. It said find the rate of change.

OpenStudy (anonymous):

26%

OpenStudy (anonymous):

265-169=96 365/96=.263013698 .26301369=26%

OpenStudy (anonymous):

is this right